OpenStudy (ledah):

5 x ^3 y^4 z divided by 10 x y^2 z^3

OpenStudy (ledah):

having a little bit of trouble with dividing this problem. not sure how to start off with division

OpenStudy (anonymous):

the equation can be re written as: 10/5 (x^3 y^4 z) (x y^2 z^3)^-1 and then you can cancel out terms, x^(3-1), y^(4-2), z^(1-3) = 2 x^2 y ^2 z^-2

OpenStudy (unklerhaukus):

\[\frac{5 x ^3 y^4 z}{10 x y^2 z^3}=\frac{5 x^2x y^2y^2 z}{2\times5 x y^2 zz^2}=\frac{\cancel{5 }x^2\cancel {x y^2}y^2 \cancel z}{2\times\cancel{5}\cancel{ x y^2} \cancel zz^2}\]